# UniGame: Turning a Unified Multimodal Model Into Its Own Adversary

### v2

@@ Abstract
Unified Multimodal Models (UMMs) have shown impressive performance in both understanding and generation with a single architecture. However, UMMs still exhibit a fundamental inconsistency: understanding favors compact embeddings, whereas generation favors reconstruction-rich representations. This structural trade-off produces misaligned decision boundaries, degraded cross-modal coherence, and heightened vulnerability under distributional and adversarial shifts. In this paper, we present UniGame, a self-adversarial post-training framework that directly targets the inconsistencies. By applying a lightweight perturber at the shared token interface, UniGame enables the generation branch to actively seek and challenge fragile understanding, turning the model itself into its own adversary. Experiments demonstrate that UniGame significantly improves the consistency ($4.6\%$). Moreover, it also achieves substantial improvements in understanding ($3.6\%$), generation (+0.02+0.02), out-of-distribution and adversarial robustness ($4.8\%$ and $6.2\%$ on NaturalBench and AdVQA). The framework is architecture-agnostic, introduces $1\%$ additional parameters, and is complementary to existing post-training methods. These results position adversarial self-play as a general and effective principle for enhancing the coherence, stability, and unified competence of future multimodal foundation models. The official code is available at: https://github.com/AIFrontierLab/UniGame.
@@ 1 Introduction
Unified Multimodal Models (UMMs) have recently demonstrated impressive capability in both visual understanding and image generation with a unified architecture [deng2025emerging, wang2024emu3, qu2025tokenflow, xie2024show, xiao2025omnigen, wu2025harmonizing]. By jointly leveraging a language model backbone and a visual tokenizer–decoder stack [li2025dual, chen2025janus], these models promise a unified interface for cross-modal reasoning, grounded perception, and controllable generation. Specifically, the large-scale pre-training establishes general multimodal capabilities, and the post-training stage (supervised fine-tuning, SFT, Figure 2(a)) can further improve their performance on downstream tasks with enhanced reliability.
Despite their great performance, UMMs exhibit structural inconsistency between their understanding and generation pathways [yan2025can, xie2025reconstruction]. This inconsistency stems from the inherently conflicting nature of the two objectives, leading to the mismatch in various aspects such as semantics (i.e., the model can answer a question correctly yet fail to generate a corresponding image, or vice versa [Gan2020, rajabi2025token]), capability (e.g., generation is harder to improve than understanding, or vice versa [radford2021learning]), and feature compactness (e.g., understanding requires more compact feature space while generation prefers oppositely). Inconsistency widely exists in real-world applications, where models frequently encounter unexpected inputs far from the training manifold, compositional combinations unseen during training, counterfactual queries, or modality conflicts such as distribution shift [Gan2020, li2024naturalbench, oh2025understanding, oh2025vittle] and adversarial attack [li2021adversarial]. If not sufficiently studied, it would greatly undermine multimodal information fusion, model robustness, and further performance improvement.
It remains challenging and unexplored to improve the consistency of UMMs, primarily due to the unclear learning objectives: only the consequences, but not the causes, are known. Therefore, recent post-training approaches tend to fill this gap using surrogate objectives. Reconstruction-based approaches (Figure 2(b)) regenerate the original images through the semantic embedding space derived from visual perception [xie2025reconstruction, yan2025can]. They optimize a unified reconstruction objective, which trains models within a closed auto-encoding loop. Reward-based methods (Figure 2(c)) typically optimize an ensembled reward function [jiang2025t2i], combining task-specific or rule-based metrics to refine the output relying on external expert models. However, both are optimized using handcrafted objectives (reconstruction or reward), which only polish model behavior on a fixed training distribution and place no explicit constraints on the two coupling branches. As a result, they reproduce behaviors within a fixed manifold rather than expanding the shared generative space, leaving the inconsistency largely unresolved.
@@
Figure 1: Qualitative and quantitative analyses of UniGame.222(a) Where consistency Score computed as the average of WISE and UnifiedBench, performance is averaged over understanding bench MMMU and generation bench like GenEval. (b) We randomly sample 100 images, extract their unified embeddings, project to 2D with UMAP [mcinnes2018umap]; colored regions visualize each method’s on-manifold coverage. (a) The performance vs. consistency score of several models, indicating significant improvement of both metrics of our models. (b) The manifold produced by SFT, reconstruction-based Post Train, and UniGame. UniGame expands the training distribution toward hard yet realistic neighborhoods.
@@
Can a UMM expose and correct its own inconsistencies from within? In this paper, motivated by the observation that adversarial signals reliably surface brittle reasoning in vision–language models [Gan2020, li2024naturalbench, oh2025understanding, rajabi2025token], we propose UniGame (Figure 3), the first self-adversarial post-training framework for UMMs. UniGame treats the generation pathway as an active adversary that searches for visually plausible, decoder-constrained perturbations that maximally challenge the understanding branch. It pushes the model beyond fixed data manifolds and produces structured adversarial samples along the uncertainty regions (Figure 1(a)). Concretely, UniGame installs a lightweight perturber at the shared visual-token interface to create bounded, structured perturbations. These perturbations are decoded into realistic adversarial images, filtered through a semantic consistency check, and stored in a hard-example buffer. The understanding branch is then optimized to correctly reason over both clean inputs and these internally generated, semantically aligned counterexamples. This forms a minimax self-play process where generation seeks to expose weaknesses while understanding learns from them, effectively expanding the shared generative manifold toward fragile yet meaningful regions (Figure 1(b); theoretical insights in Appendix G.3). Compared to existing efforts, UniGame explicitly converts representational weaknesses into decoded, semantically coherent counterexamples that efficiently harden understanding (Figure 2). Empirically, UniGame uncovers richer, actionable failure modes (e.g., counting, fine attributes, and occlusion in Figure 5(a)) and improves the consistency, performance, and robustness.
@@
    Empirical improvement. UniGame yields improvements in consistency (4.6%), understanding (+3.6%), generation (+0.02), OOD (+4.8%) and adversarial robustness (+6.2%).
@@ 2 Related Work
Unified Multimodal Models. UMMs aim to combine multimodal understanding and generation within a single backbone, enabling compact deployment and richer cross-modal reasoning [chen2025blip3, wang2024emu3, team2024chameleon, chen2025janus, qu2025tokenflow, deng2025emerging, qu2025tokenflow]. Among these, BLIP3-o [chen2025blip3] explores hybrid autoregressive and diffusion training recipes to balance understanding and generation fidelity. Emu3 [wang2024emu3] treats images and text as an interleaved token stream and scales next-token prediction to unified multimodal outputs. TokenFlow [qu2025tokenflow] focuses on the tokenizer layer and introduces dual-granularity codebooks to reconcile the conflicting demands of discriminative understanding vs reconstructive generation. Despite architectural innovation, UMMs continue to face the inconsistency issue: the representation granularity and objective tension that underlie understanding and generation still cause ambiguous shared embeddings and latent failure modes that remain insufficiently addressed.
Post-training of UMMs. Aiming to resolve the inconsistency issue, existing post-training methods can be categorized into these types: (i) reconstruction and semantic-alignment losses to encourage fidelity [xie2025reconstruction, yan2025can]; (ii) RL/reward-based optimization to directly optimize downstream metrics [Ramesh2024, jiang2025t2i]. Each family improves either fidelity or robustness [Li2023, wu2024janus, xie2025reconstruction, Gan2020, rajabi2025token, Ramesh2024, qu2025tokenflow], For instance, RecA [xie2025reconstruction] leverages reconstruction alignment by conditioning generation on understanding embeddings and using reconstruction losses to bring representations closer, while T2I-R1 [jiang2025t2i] enhances image generation through collaborative semantic- and token-level chain-of-thought combined with reinforcement learning. In the vision–language domain, AT has shown potential: VILLA introduces large-scale embedding-space perturbations across image and text modalities, improving robustness and generalization [Gan2020, rajabi2025token]. Nevertheless, most methods either improve alignment or generation separately, and adversarial mechanisms are rarely incorporated into full UMMs. They commonly fail in one crucial respect: they do not exploit generation as an active adversarial process to strengthen the understanding branch. Specifically, adversarial or unguided embedding perturbations often produce off-manifold samples—unrealistic or semantically invalid artifacts, while reconstruction objectives do not intentionally surface decision-critical failure modes. Although RL-based schemes are effective, they are computationally costly and do not guarantee that these discovered examples remain decodable and semantically plausible [Ramesh2024, jiang2025t2i]. This leaves a key gap in truly improving the consistency of the understanding and generation pathways.
(a) Supervised fine-tuning
(b) Reconstruction-base approaches
(c) Reward-based approaches
(d) Ours: Minmax optimization
Figure 2: Illustration of four different post-training paradigms.
@@ 3.2 Motivation
Existing efforts [xie2025reconstruction, jiang2025t2i] optimize for individual goals on a fixed set of data distributions, and place no explicit constraints on the two coupling training objectives. They encourage cooperative training, reproducing existing samples rather than expanding coverage of the shared manifold, where boundary behavior is most fragile, thus aggravating inconsistency. We argue that improving consistency requires expanding this shared manifold, especially around decision boundaries, instead of merely polishing the model within its comfort regions.
Considering the unified architecture of UMMs: can we improve the consistency within the model itself? We turn to adversarial training [Madry2017], which creates adversarial perturbations to explore understanding failures [Gan2020, rajabi2025token]. This indicates that adversarial signals, if properly constrained, can serve as an effective mechanism for regularizing the decision boundaries of UMMs. Within the UMM architecture, we focus on converting generative priors into decodable adversarial cases that (i) remain semantically valid and (ii) reliably expose genuine reasoning failures in the understanding branch. This intuitively motivates the self-play training paradigm that makes best use of both of the understanding and generation branches as a minimax optimization framework. The generative pathway no longer passively follows alignment objectives: it is explicitly trained to produce realistic, on-manifold adversarial cases that challenge the understanding module, while the understanding branch is optimized to solve these internally generated challenges.
@@ 3.3 Overview of UniGame
  • Hard-sample buffer $\mathcal{B}$: A component that scores decoded candidates and stores hard, semantically plausible examples for replay via semantic-consistency check [radford2021learning]:
@@
We refer the original generation branch as the clean path, then the perturber introduces a perturbed path. In UniGame, a frozen visual encoder maps an input image to visual tokens at the unified interface. In the clean path, it is forwarded to the understanding head for supervised training with the query $q$; In the perturbed path, the embeddings with perturbations generated by $C$ are decoded by $G$ into candidate images. The semantically consistent candidates are stored in the buffer $\mathcal{B}$. During training, the understanding module learns from hard examples from $\mathcal{B}$ and clean samples, and the perturber $C$ is optimized to generate challenging yet plausible cases. The vision encoder (SigLIP [zhai2023sigmoid]) is frozen, and only LoRA [hu2022lora] adapters on the LLM backbone and the Perturber $C$ are trainable.
@@ 3.4 The Self-Play Training Process
$\begin{split}\mathcal{L}_{U}(\theta_{U})=\;\mathbb{E}_{\mathrm{clean}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big]+\beta\,\mathbb{E}_{\mathcal{B}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big],\end{split}$ (6)
@@
$\begin{split}\mathcal{L}_{C}(\theta_{C};\theta_{U})=\;\mathbb{E}_{\mathrm{clean}}\mathrm{CE}\big(p_{U}(\hat{a}\mid\mathrm{Enc}(G(C(\hat{\mathbf{z}};\theta_{C}))),q;\theta_{U}),a\big)-\lambda\|\boldsymbol{\delta}\|^{2}.\end{split}$ (7)
@@ 3.5 Discussion
UniGame vs. GANs. UniGame is different from conventional GANs [Goodfellow2014gan] that train a generator to fool a discriminator [Goodfellow2014gan, Radford2015]. First, GAN needs an extra discriminator for the adversarial game, while ours can operate within a UMM to leverage its own understanding and generation branches. Second, GANs primarily focus on generation tasks while ours targets both generation and understanding, involving more complex training and optimization process.
UniGame vs. Adversarial Training (AT). UniGame is the first attempt of applying AT [Madry2017, Goodfellow2014adv] to UMMs, but has the following differences. First, AT is mainly used for enhancing adversarial robustness, while ours explores AT for consistency improvement. Second, UniGame differs fundamentally from prior AT by enforcing decoder-constrained, on-manifold image perturbations, enabling self-generated adversarial cases that remain semantically meaningful.
Extensibility. UniGame is agnostic to most UMM architectures and post-training approaches. First, since it is a general training framework that only introduces a lightweight trainable perturber, it is flexible to be integrated into most UMMs. Second, it does not conflict with existing methods, but can serve as their complement for further improvement on consistency and performance. We further explore the intergration of UniGame with emerging post-training method e.g., [jiang2025t2i, xie2025reconstruction], we train their post-trained model and demonstrate further improvements (see §4.6).
@@ 4 Experiments
In this section, we conducted extensive experiments to evaluate the proposed UniGame.
@@ 4.1 Experimental Setup
Tasks and datasets. We evaluated UniGame on popular benchmarks. Specifically, VQAv2 [balanced_vqa_v2], MMMU [yue2024mmmu], POPE [li2023evaluating], and MMBench  [liu2024mmbench] are adopted for understanding evaluation and GenEval [ghosh2023geneval] is employed for generation evaluation. For the evaluation of consistency, we report WISE score [niu2025wise] and UnifiedBench [yan2025can]. For OOD robustness, we adopt NaturalBench [li2024naturalbench], a challenging benchmark comprising real-world images captured in natural, uncontrolled environments (e.g., low lighting, occlusion, unusual viewpoints) that test robust visual reasoning. For adversarial robustness, we adopt AdVQA [li2021adversarial], an adversarially constructed VQA dataset where questions are intentionally designed to mislead models through linguistic ambiguity and visual distractors.
Implementation details. We implemented UniGame on a popular Janus-Pro-7B [chen2025janus] UMM for main experiments, and further validated with two toy models that simulate distinct UMMs designs, similar to UAE [yan2025can]. The perturber is implemented as a 3-layer MLP operating on the shared visual-token space $\mathbb{R}^{N\times H}$, adding only 2.12.1M parameters and outputting token-wise perturbations followed by normalization and clipping. We adopt the open-source VQAv2 training set and CC3M [sharma2018conceptual]. More training details are in Appendix B.
Table 1: Consistency [^6] Consistency Score = $0.6\times\text{UnifiedBench}+0.4\times(\text{WISE}\times 100)$, jointly assessing self-consistency in understanding generated content (UnifiedBench) and prompt-image alignment (WISE), results marked with show improvement over base model. evaluation on UnifiedBench and WISE ($\uparrow$). Models marked with denote our base model.
@@
BAGEL [deng2025emerging] 14B $83.48$ $0.41$ $41.95$ $66.49$
UniWorld-V1 [lin2025uniworld] 12B $78.99$ $0.35$ $39.67$ $61.39$
BLIP-3o [chen2025blip3] 8B $76.56$ $0.39$ $38.48$ $61.54$
OmniGen2 [xiao2025omnigen] 7B $83.31$ $0.30$ $41.81$ $61.99$
Janus-Pro [chen2025janus] 7B $82.77$ $0.35$ $41.54$ $63.66$
Harmon [wu2025harmonizing] 1.5B $65.41$ $0.41$ $32.90$ $55.65$
Show-o [xie2024show] 1.3B $69.16$ $0.30$ $34.73$ $53.50$
Janus-Pro+SFT [chen2025janus] 7B $83.20$ $0.37$ $41.79$ 64.72(+1.06)64.72~(+1.06)
Harmon+RecA [xie2025reconstruction] 1.5B $66.94$ $0.40$ $33.67$ 56.16(+0.51)56.16~(+0.51)
@@
Baselines. We compare UniGame to two categories of baselines: (1) Different UMMs: (i) Auto-regressive models [chen2025janus, team2024chameleon, qu2025tokenflow, ge2024seed, xie2024show], which unify understanding and generation through next-token prediction in a shared token space, with improvements in vision tokenizers; (ii) Diffusion-based models [chen2025blip3, xiao2025omnigen], which leverage latent diffusion for generation while maintaining autoregressive understanding (e.g., BLIP-3o [chen2025blip3], OmniGen2 [xiao2025omnigen]); (iii) Hybrid architectures [lin2025uniworld, deng2025emerging], which employ specialized modules for different modalities (e.g., UniWorld-V1 [lin2025uniworld], BAGEL [deng2025emerging]). (2) Post-training methods: (i) Reconstruction-based alignment, which uses caption-then-reconstruct cycles to enhance understanding-generation consistency (e.g., RecA [xie2025reconstruction]); (ii) Reward-based approaches, which use reward function to refine their outputs(e.g, [jiang2025t2i]) Unlike these methods, UniGame introduces decoder-constrained self-adversarial training, converting latent inconsistencies into visually coherent counterexamples to improve reasoning robustness while preserving generation fidelity.

### v3

@@ Abstract
Unified Multimodal Models (UMMs) have shown impressive performance in both understanding and generation with a single architecture. However, UMMs still exhibit a fundamental inconsistency: understanding favors compact embeddings, whereas generation favors reconstruction-rich representations. This structural trade-off produces misaligned decision boundaries, degraded cross-modal coherence, and heightened vulnerability under distributional and adversarial shifts. In this paper, we present UniGame, a self-adversarial post-training framework that directly targets the inconsistencies. By applying a lightweight perturber at the shared token interface, UniGame enables the generation branch to actively seek and challenge fragile understanding, turning the model itself into its own adversary. Experiments demonstrate that UniGame significantly improves the consistency ($4.6\%$). Moreover, it also achieves substantial improvements in understanding ($3.6\%$), generation (+0.02+0.02) on GenEval, out-of-distribution and adversarial robustness ($4.8\%$ and $6.2\%$ on NaturalBench and AdVQA). The framework is architecture-agnostic, introduces $1\%$ additional parameters, and is complementary to existing post-training methods. These results position adversarial self-play as a general and effective principle for enhancing the coherence, stability, and unified competence of future multimodal models. Code is available at https://github.com/AIFrontierLab/TorchUMM.
@@ 1 Introduction
Unified Multimodal Models (UMMs) have recently demonstrated impressive capability in both visual understanding and image generation with a unified architecture [4, 37, 28, 42, 41, 40]. By jointly leveraging a language model backbone and a visual tokenizer–decoder stack [18, 3], these models promise a unified interface for cross-modal reasoning, grounded perception, and controllable generation. Specifically, the large-scale pre-training establishes general multimodal capabilities, and the post-training stage (supervised fine-tuning, SFT, Figure 2(a)) can further improve their performance on downstream tasks with enhanced reliability.
@@
Figure 1: Qualitative and quantitative analyses of UniGame.222(a) Where consistency Score computed as the average of WISE and UnifiedBench, performance is averaged over understanding bench MMMU and generation bench like GenEval. (b) We randomly sample 100 images, extract their unified embeddings, project to 2D with UMAP [22]; colored regions visualize each method’s on-manifold coverage. (a) The performance vs. consistency score of several models, indicating significant improvement of both metrics of our models. (b) The manifold produced by SFT, reconstruction-based Post Train, and UniGame. UniGame expands the training distribution toward hard yet realistic neighborhoods.
(a) Supervised fine-tuning
(b) Reconstruction-base approaches
(c) Reward-based approaches
(d) Ours: Minmax optimization
Figure 2: Illustration of four different post-training paradigms.
Despite their great performance, UMMs exhibit structural inconsistency between their understanding and generation pathways [43, 38]. This inconsistency stems from the inherently conflicting nature of the two objectives, leading to the mismatch in various aspects such as semantics (i.e., the model can answer a question correctly yet fail to generate a corresponding image, or vice versa [6, 31]), capability (e.g., generation is harder to improve than understanding, or vice versa [29]), and feature compactness (e.g., understanding requires more compact feature space while generation prefers oppositely). Inconsistency widely exists in real-world applications, where models frequently encounter unexpected inputs far from the training manifold, compositional combinations unseen during training, counterfactual queries, or modality conflicts such as distribution shift [6, 14, 24, 25] and adversarial attack [16]. If not sufficiently studied, it would greatly undermine multimodal information fusion, model robustness, and further performance improvement.
It remains challenging and unexplored to improve the consistency of UMMs, primarily due to the unclear learning objectives: only the consequences, but not the causes, are known. Therefore, recent post-training approaches tend to fill this gap using surrogate objectives. Reconstruction-based approaches (Figure 2(b)) regenerate the original images through the semantic embedding space derived from visual perception [38, 43]. They optimize a unified reconstruction objective, which trains models within a closed auto-encoding loop. Reward-based methods (Figure 2(c)) typically optimize an ensembled reward function [13], combining task-specific or rule-based metrics to refine the output relying on external expert models. However, both are optimized using handcrafted objectives (reconstruction or reward), which only polish model behavior on a fixed training distribution and place no explicit constraints on the two coupling branches. As a result, they reproduce behaviors within a fixed manifold rather than expanding the shared generative space, leaving the inconsistency largely unresolved.
@@
Can a UMM expose and correct its own inconsistencies from within? In this paper, motivated by the observation that adversarial signals reliably surface brittle reasoning in vision–language models [6, 14, 24, 31], we propose UniGame (Figure 3), the first self-adversarial post-training framework for UMMs. UniGame treats the generation pathway as an active adversary that searches for visually plausible, decoder-constrained perturbations that maximally challenge the understanding branch. It pushes the model beyond fixed data manifolds and produces structured adversarial samples along the uncertainty regions (Figure 1(a)). Concretely, UniGame installs a lightweight perturber at the shared visual-token interface to create bounded, structured perturbations. These perturbations are decoded into realistic adversarial images, filtered through a semantic consistency check, and stored in a hard-example buffer. The understanding branch is then optimized to correctly reason over both clean inputs and these internally generated, semantically aligned counterexamples. This forms a minimax self-play process where generation seeks to expose weaknesses while understanding learns from them, effectively expanding the shared generative manifold toward fragile yet meaningful regions (Figure 1(b); theoretical insights in Appendix G.3). Compared to existing efforts, UniGame explicitly converts representational weaknesses into decoded, semantically coherent counterexamples that efficiently harden understanding (Figure 2). Empirically, UniGame uncovers richer, actionable failure modes (e.g., counting, fine attributes, and occlusion in Figure 5(a)) and improves the consistency, performance, and robustness.
@@
    Empirical improvement. UniGame yields improvements in consistency (4.6%), understanding (+3.6%), generation (+0.02 on GenEval), OOD (+4.8%) and adversarial robustness (+6.2%).
@@ 2 Related Work
Unified Multimodal Models. UMMs aim to combine multimodal understanding and generation within a single backbone, enabling compact deployment and richer cross-modal reasoning [1, 37, 36, 2, 28, 4, 28]. Among these, BLIP3-o [1] explores hybrid autoregressive and diffusion training recipes to balance understanding and generation fidelity. Emu3 [37] treats images and text as an interleaved token stream and scales next-token prediction to unified multimodal outputs. TokenFlow [28] focuses on the tokenizer layer and introduces dual-granularity codebooks to reconcile the conflicting demands of discriminative understanding vs reconstructive generation. Despite architectural innovation, UMMs continue to face the inconsistency issue: the representation granularity and objective tension that underlie understanding and generation still cause ambiguous shared embeddings and latent failure modes that remain insufficiently addressed.
Post-training of UMMs. Aiming to resolve the inconsistency issue, existing post-training methods can be categorized into these types: (i) reconstruction and semantic-alignment losses to encourage fidelity [38, 43]; (ii) RL/reward-based optimization to directly optimize downstream metrics [32, 13]. Each family improves either fidelity or robustness [15, 3, 38, 6, 31, 32, 28], For instance, RecA [38] leverages reconstruction alignment by conditioning generation on understanding embeddings and using reconstruction losses to bring representations closer, while T2I-R1 [13] enhances image generation through collaborative semantic- and token-level chain-of-thought combined with reinforcement learning. In the vision–language domain, AT has shown potential: VILLA introduces large-scale embedding-space perturbations across image and text modalities, improving robustness and generalization [6, 31]. Nevertheless, most methods either improve alignment or generation separately, and adversarial mechanisms are rarely incorporated into full UMMs. They commonly fail in one crucial respect: they do not exploit generation as an active adversarial process to strengthen the understanding branch. Specifically, adversarial or unguided embedding perturbations often produce off-manifold samples—unrealistic or semantically invalid artifacts, while reconstruction objectives do not intentionally surface decision-critical failure modes. Although RL-based schemes are effective, they are computationally costly and do not guarantee that these discovered examples remain decodable and semantically plausible [32, 13]. This leaves a key gap in truly improving the consistency of the understanding and generation pathways.
@@ 3.2 Motivation
Existing efforts [38, 13] optimize for individual goals on a fixed set of data distributions, and place no explicit constraints on the two coupling training objectives. They encourage cooperative training, reproducing existing samples rather than expanding coverage of the shared manifold, where boundary behavior is most fragile, thus aggravating inconsistency. We argue that improving consistency requires expanding this shared manifold, especially around decision boundaries, instead of merely polishing the model within its comfort regions.
Considering the unified architecture of UMMs: can we improve the consistency within the model itself? We turn to adversarial training [21], which creates adversarial perturbations to explore understanding failures [6, 31]. This indicates that adversarial signals, if properly constrained, can serve as an effective mechanism for regularizing the decision boundaries of UMMs. Within the UMM architecture, we focus on converting generative priors into decodable adversarial cases that (i) remain semantically valid and (ii) reliably expose genuine reasoning failures in the understanding branch. This intuitively motivates the self-play training paradigm that makes best use of both of the understanding and generation branches as a minimax optimization framework. The generative pathway no longer passively follows alignment objectives: it is explicitly trained to produce realistic, on-manifold adversarial cases that challenge the understanding module, while the understanding branch is optimized to solve these internally generated challenges.
@@ 3.3 Overview of UniGame
  • Hard-sample buffer $\mathcal{B}$: A component that scores decoded candidates and stores hard, semantically plausible examples for replay via semantic-consistency check [29]:
@@
We refer the original generation branch as the clean path, then the perturber introduces a perturbed path. In UniGame, a frozen visual encoder maps an input image to visual tokens at the unified interface. In the clean path, it is forwarded to the understanding head for supervised training with the query $q$; In the perturbed path, the embeddings with perturbations generated by $C$ are decoded by $G$ into candidate images. The semantically consistent candidates are stored in the buffer $\mathcal{B}$. During training, the understanding module learns from hard examples from $\mathcal{B}$ and clean samples, and the perturber $C$ is optimized to generate challenging yet plausible cases. The vision encoder (SigLIP [46]) is frozen, and only LoRA [12] adapters on the LLM backbone and the Perturber $C$ are trainable.
@@ 3.4 The Self-Play Training Process
$\begin{split}\mathcal{L}_{U}(\theta_{U})=\;&\mathbb{E}_{\mathrm{clean}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big]\\ &+\beta\,\mathbb{E}_{\mathcal{B}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big],\end{split}$ (6)
@@
[…]
@@ 3.5 Discussion
UniGame vs. GANs. UniGame is different from conventional GANs [9] that train a generator to fool a discriminator [9, 30]. First, GAN needs an extra discriminator for the adversarial game, while ours can operate within a UMM to leverage its own understanding and generation branches. Second, GANs primarily focus on generation tasks while ours targets both generation and understanding, involving more complex training and optimization process.
UniGame vs. Adversarial Training (AT). UniGame is the first attempt of applying AT [21, 10] to UMMs, but has the following differences. First, AT is mainly used for enhancing adversarial robustness, while ours explores AT for consistency improvement. Second, UniGame differs fundamentally from prior AT by enforcing decoder-constrained, on-manifold image perturbations, enabling self-generated adversarial cases that remain semantically meaningful.
Extensibility. UniGame is agnostic to most UMM architectures and post-training approaches. First, since it is a general training framework that only introduces a lightweight trainable perturber, it is flexible to be integrated into most UMMs. Second, it does not conflict with existing methods, but can serve as their complement for further improvement on consistency and performance. We further explore the intergration of UniGame with emerging post-training method e.g., [13, 38], we train their post-trained model and demonstrate further improvements (see §4.6).
@@ 4.1 Experimental Setup
Tasks and datasets. We evaluated UniGame on popular benchmarks. Specifically, VQAv2 [11], MMMU [45], POPE [17], and MMBench  [20] are adopted for understanding evaluation and GenEval [8] is employed for generation evaluation. For the evaluation of consistency, we report WISE score [23] and UnifiedBench [43]. For OOD robustness, we adopt NaturalBench [14], a challenging benchmark comprising real-world images captured in natural, uncontrolled environments (e.g., low lighting, occlusion, unusual viewpoints) that test robust visual reasoning. For adversarial robustness, we adopt AdVQA [16], an adversarially constructed VQA dataset where questions are intentionally designed to mislead models through linguistic ambiguity and visual distractors.
Implementation details. We implemented UniGame on a popular Janus-Pro-7B [2] UMM for main experiments, and further validated with two toy models that simulate distinct UMMs designs, similar to UAE [43]. The perturber is implemented as a 3-layer MLP operating on the shared visual-token space $\mathbb{R}^{N\times H}$, adding only 2.12.1M parameters and outputting token-wise perturbations followed by normalization and clipping. We adopt the open-source VQAv2 training set and CC3M [35]. More training details are in Appendix B.
Table 1: Consistency [^6] Consistency Score = $0.6\times\text{UnifiedBench}+0.4\times(\text{WISE}\times 100)$, jointly assessing self-consistency in understanding generated content (UnifiedBench) and prompt-image alignment (WISE), the $0.6/0.4$ weighting reflects the evaluation data ratio. show improvement over base model.evaluation on UnifiedBench and WISE ($\uparrow$). Models marked with denote our base model.
@@
BAGEL [4] 14B $83.48$ $0.41$ $41.95$ $66.49$
UniWorld-V1 [19] 12B $78.99$ $0.35$ $39.67$ $61.39$
BLIP-3o [1] 8B $76.56$ $0.39$ $38.48$ $61.54$
OmniGen2 [41] 7B $83.31$ $0.30$ $41.81$ $61.99$
Janus-Pro [2] 7B $82.77$ $0.35$ $41.54$ $63.66$
Harmon [40] 1.5B $65.41$ $0.41$ $32.90$ $55.65$
Show-o [42] 1.3B $69.16$ $0.30$ $34.73$ $53.50$
Janus-Pro+SFT [2] 7B $83.20$ $0.37$ $41.79$ 64.72(+1.06)64.72~(+1.06)
Harmon+RecA [38] 1.5B $66.94$ $0.40$ $33.67$ 56.16(+0.51)56.16~(+0.51)
@@
Baselines. We compare UniGame to two categories of baselines: (1) Different UMMs: (i) Auto-regressive models [2, 36, 28, 7, 42], which unify understanding and generation through next-token prediction in a shared token space, with improvements in vision tokenizers; (ii) Diffusion-based models [1, 41], which leverage latent diffusion for generation while maintaining autoregressive understanding (e.g., BLIP-3o [1], OmniGen2 [41]); (iii) Hybrid architectures [19, 4], which employ specialized modules for different modalities (e.g., UniWorld-V1 [19], BAGEL [4]). (2) Post-training methods: (i) Reconstruction-based alignment, which uses caption-then-reconstruct cycles to enhance understanding-generation consistency (e.g., RecA [38]); (ii) Reward-based approaches, which use reward function to refine their outputs(e.g, [13]) Unlike these methods, UniGame introduces decoder-constrained self-adversarial training, converting latent inconsistencies into visually coherent counterexamples to improve reasoning robustness while preserving generation fidelity.